OpenStudy (anonymous):

what is the equation of the line parallel to the line y = 5x + 2 and passes through the point (1, -3)?

OpenStudy (anonymous):

what is the slope of \(y=5x+2\) ?

OpenStudy (anonymous):

Satellite got this one I will back down.

OpenStudy (anonymous):

2

OpenStudy (dumbcow):

y = mx +b slope = m

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

\[\huge y=\color{blue} mx+\color{red}b\] \(\color{blue}m\) is the slope and \(\color{red}b\) is the \(y\) intercept

OpenStudy (anonymous):

So it's 5

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

and so a parallel line will also have slope \(5\) your job is to find the line with slope \(5\) through the point \((1,-3)\) do you know how to do that?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok you want to write the point slope formula \[y-y_1=m(x-x_1)\] with \[m=5,x_1=1,y_1=-3\]

OpenStudy (anonymous):

do you understand what i am asking you to write?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Y -3 = 5 ( x - 1)

OpenStudy (anonymous):

close, but be careful \(y_1=-3\) so \(y-y_1=y-(-3)=y+3\)

OpenStudy (anonymous):

you should have \[y+3=5(x-1)\] and now multiply out on the right using the distributive law

OpenStudy (anonymous):

Y+3=5x-5

OpenStudy (anonymous):

right then subtract \(3\) from both sides to put it in the form you are asked for

OpenStudy (anonymous):

Y=5x-8

OpenStudy (anonymous):

Is that correct answer?

OpenStudy (anonymous):

@Alisonflowers2015 This is pretty simple the slopes of two parallel line are always equal s=3o according to the st. line equation y=mx+c where m is your slope the slope of first line y=5x+2 is 5 i.e m1=5 Since second line in parallel to the first line its slope m2=5 Now you cannot use y=mx+c to find the equation of second line because you do not know the y-intercept i.e c. So now you would have to use point slope form of equation to find the equation of the second line since you have the co-ordinates point slope form of st. line equation (y-y1)=m(x-x1) taking y1=-3 and x1=1 (y+3)=5(x-1) y=5x-5-3 y=5x-8 y=5x-8 is the equation of the line which is parallel to line y=5x+2